OpenStudy (anonymous):

Simplify. The variable represents a nonnegative real number. √147b^3/b

OpenStudy (anonymous):

\[\sqrt{147b ^{3}}/b=\sqrt{49*3*b ^{2}*b}/b=7\sqrt{3b}\]